OpenStudy (anonymous):

Find the range of values of x for (x-3)^2 < or equal 14 - 2x

OpenStudy (anonymous):

\[x \in [-1,5]\]

OpenStudy (anonymous):

\[(x - 3)^2 \le 14 - 2x\] \[(x -3)^2 -14 +2x \le 0\] \[\text{Let } f(x) = (x -3)^2 -14 +2x\] \[f(x) \le 0 \] \[(x -3)^2 -14 +2x \le 0 \implies x^2 + 9 - 6x + 2x -14 \le 0 \implies x^2 - 4x - 5 \le 0 \] \[\implies x^2 - 5x + x - 5 \le 0 \implies x(x -5) + 1(x -5) \le 0 \implies (x - 5)(x +1)\le 0 \] \( ++++\) \( ---- \) \( ++++\) <-----------------------------|-------------------------|-------------------------------------------------> \(-\infty\) -1 5 \(\infty\) Hence, \(x \in [-1,5]\) is the Solution.

OpenStudy (anonymous):

That's Cheating :-p lol

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I was kidding

OpenStudy (anonymous):

i answered almost 10 minutes before u :P

OpenStudy (anonymous):

sorry, 4 minutes..

OpenStudy (anonymous):

3 and a Half :-P but I have the solution too

OpenStudy (anonymous):

great :D

OpenStudy (anonymous):

nobody's giving me a medal :(

OpenStudy (anonymous):

You Just Got one! :-D